# Model-assisted inference for dynamic causal effects in staggered rollout cluster randomized experiments

### v2

@@ 3.2.1 Without covariates
Then, the CR variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{I}}\bm{Q}(a,a^{\prime})^{T}$, and the CR variance estimator for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ is
@@
Under Assumptions 1 and 2, and regularity conditions (C1) and (C2) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{j,\mathrm{I}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if we further assume $\max_{i}\pi_{ij\cdot}=o(I^{-2/3})$ for $j=1,\ldots,J$ and $\bm{V}(\widetilde{\bm{\epsilon}}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\widetilde{\bm{\epsilon}}_{i\cdot})(a,a^{\prime}))$, where $\widetilde{\bm{\epsilon}}_{i\cdot}=(\widetilde{\epsilon}_{i1\cdot}(a),\ldots,\widetilde{\epsilon}_{iJ\cdot}(a))^{T}$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}=\bm{V}_{c}(\widetilde{\bm{\epsilon}}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
@@ 3.2.2 With covariates
where the centered individual-level covariate vector $\bm{X}_{ijk}^{c}=\bm{X}_{ijk}-\overline{\bm{X}}_{\cdot j\cdot}$ with $\overline{\bm{X}}_{\cdot j\cdot}=\sum_{i=1}^{I}\pi_{ij\cdot}\overline{\bm{X}}_{ij\cdot}$ and $\overline{\bm{X}}_{ij\cdot}=\sum_{k=1}^{N_{ij}}\pi_{ijk}\bm{X}_{ijk}\allowbreak/\pi_{ij\cdot}$, and $\bm{\gamma}_{j}(a)$ is the adoption-time-specific parameter vector. Similar to $\widehat{\beta}_{j,\mathrm{I}}(a)$, with individual weight $\pi_{ijk}$, we can obtain WLS estimates $\widehat{\beta}_{j,\mathrm{I}}^{\mathrm{adj}}(a)$ and $\widehat{\bm{\gamma}}_{j,\mathrm{I}}(a)$ via fitting (7). A covariate-adjusted estimator for $\tau_{j}(a,a^{\prime})$ can be obtained in the form of $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})=\widehat{\beta}_{j,\mathrm{I}}^{\mathrm{adj}}(a)-\widehat{\beta}_{j,\mathrm{I}}^{\mathrm{adj}}(a^{\prime})$, with $\widehat{\beta}_{j,\mathrm{I}}^{\mathrm{adj}}(a)=\overline{y}_{\cdot j\cdot}(a)-\overline{\bm{x}}_{\cdot j\cdot}^{c}(a)\widehat{\bm{\gamma}}_{j,\mathrm{I}}(a)$ where $\overline{\bm{x}}_{\cdot j\cdot}^{c}(a)=\sum_{i=1}^{I}\pi_{ij\cdot}G_{i}(a)\overline{\bm{X}}_{ij\cdot}^{c}/\allowbreak\pi_{\cdot j\cdot}(a)$ and $\overline{\bm{X}}_{ij\cdot}^{c}=\sum_{k=1}^{N_{ij}}\pi_{ijk}\bm{X}_{ijk}^{c}/\pi_{ij\cdot}$. The covariate-adjusted estimator $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})$ can be viewed as the multi-arm clustered extension of the regression estimator using a covariate-treatment indicator interaction model in Lin (2013).
@@
$\bm{E}_{\mathrm{I}}^{\mathrm{adj}}$ $=\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}\right)^{-1}\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{I}}\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{adj}}(\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}\right)$
$\qquad\times\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}})^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{adj}}$ is the vector of residuals from the WLS fit of (7), $\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}$ is of dimensions $N_{i\cdot}\times J(1+p_{x})(J+1)$ with $\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}=((\bm{D}_{i1,\mathrm{I}}^{\mathrm{adj}})^{T}\cdots(\bm{D}_{iJ,\mathrm{I}}^{\mathrm{adj}})^{T})^{T}$ and the $k$-th row of $\bm{D}_{ij,\mathrm{I}}^{\mathrm{adj}}$ being $(\bm{G}_{i}(\mathcal{A})^{T}\otimes\overrightarrow{\mathbbm{1}}_{ijk}^{T},\bm{G}_{i}(\mathcal{A})^{T}\otimes(\overrightarrow{\mathbbm{1}}_{ijk}^{T}\otimes\bm{X}_{ijk}^{c}))$. The CR variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}(a,a^{\prime})^{T}$, where $\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}$ is the submatrix consisting of the first $J(J+1)$ columns and $J(J+1)$ rows of $\bm{E}_{\mathrm{I}}^{\mathrm{adj}}$. Let $\bm{\gamma}_{j,\mathrm{I}}(a)$ denote the WLS coefficient vectors obtained using individual-level data if the full set of potential outcomes is available. That is, regress $Y_{ijk}(a)$ on $(1,\bm{X}_{ijk}^{c})$ with weight $\pi_{ijk}$ assuming $Y_{ijk}(a)$ is available for each individual $k$ from cluster $i$ in period $j$. We further define the scaled cluster-period totals for non-centered and centered covariates as $\widetilde{\bm{X}}_{ij\cdot}=I\pi_{ij\cdot}\overline{\bm{X}}_{ij\cdot}$ and $\widetilde{\bm{X}}_{ij\cdot}^{c}=I\pi_{ij\cdot}\overline{\bm{X}}_{ij\cdot}^{c}$, respectively. The theoretical properties of $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})$ and $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}$ are given in Theorem 2.
@@
Define $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{T}$. Under Assumptions 1 and 2, and regularity conditions (C1) - (C3) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $\max_{i}\pi_{ij\cdot}=o(I^{-2/3})$ for $j=1,\ldots,J$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
@@
$\bm{E}_{\mathrm{I}}^{\mathrm{ancova}}$ $=\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}})^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}\right)^{-1}\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}})^{T}\bm{\Pi}_{i,\mathrm{I}}\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{ancova}}(\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{ancova}})^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}\right)$
$\qquad\times\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}})^{T}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{ancova}}$ is the vector of residuals from the WLS fit of (8), $\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}$ is of dimensions $N_{i\cdot}\times J(J+1+p_{x})$ with $\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}=((\bm{D}_{i1,\mathrm{I}}^{\mathrm{ancova}})^{T}\cdots(\bm{D}_{iJ,\mathrm{I}}^{\mathrm{ancova}})^{T})^{T}$ and the $k$-th row of $\bm{D}_{ij,\mathrm{I}}^{\mathrm{ancova}}$ being $(\bm{G}_{i}(\mathcal{A})^{T}\otimes\overrightarrow{\mathbbm{1}}_{ijk}^{T},\overrightarrow{\mathbbm{1}}_{ijk}^{T}\otimes\bm{X}_{ijk}^{c})$. The CR variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{ancova}}\bm{Q}(a,a^{\prime})^{T}$. The theoretical properties of $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})$ and $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})\}$ are given in the following Corollary 1.
@@
Define $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-\sum_{a\in\mathcal{A}}q(a)\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\allowbreak\ldots,r_{iJ\cdot}(a))^{T}$. Under Assumptions 1 and 2, and regularity conditions (C1) - (C3) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $\max_{i}\pi_{ij\cdot}=o(I^{-2/3})$ for $j=1,\ldots,J$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\allowbreak\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})\}=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.

### v3

@@ 3.2.1 Without covariates
Then, the CR variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{I}}\bm{Q}(a,a^{\prime})^{\top}$, and the CR variance estimator for $\mathrm{se}^{2}\{\widehat{\tau}_{j,\mathrm{I}}(a,a^{\prime})\}$ is
@@
Under Assumptions 1 and 2, and regularity conditions (C1) and (C2) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{j,\mathrm{I}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if we further assume $\max_{i}\pi_{ij\cdot}=o(I^{-2/3})$ for $j=1,\ldots,J$ and $\bm{V}(\widetilde{\bm{\epsilon}}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\widetilde{\bm{\epsilon}}_{i\cdot})(a,a^{\prime}))$, where $\widetilde{\bm{\epsilon}}_{i\cdot}=(\widetilde{\epsilon}_{i1\cdot}(a),\ldots,\widetilde{\epsilon}_{iJ\cdot}(a))^{\top}$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}(a,a^{\prime})\}=\bm{V}_{c}(\widetilde{\bm{\epsilon}}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
@@ 3.2.2 With covariates
where the centered individual-level covariate vector $\bm{X}_{ijk}^{c}=\bm{X}_{ijk}-\overline{\bm{X}}_{\cdot j\cdot}$ with $\overline{\bm{X}}_{\cdot j\cdot}=\sum_{i=1}^{I}\pi_{ij\cdot}\overline{\bm{X}}_{ij\cdot}$ and $\overline{\bm{X}}_{ij\cdot}=\sum_{k=1}^{N_{ij}}\pi_{ijk}\bm{X}_{ijk}\allowbreak/\pi_{ij\cdot}$, and $\bm{\gamma}_{j}(a)$ is the adoption-time-specific parameter vector. Similar to $\widehat{\beta}_{j,\mathrm{I}}(a)$, with individual weight $\pi_{ijk}$, we can obtain WLS estimates $\widehat{\beta}_{j,\mathrm{I}}^{\mathrm{adj}}(a)$ and $\widehat{\bm{\gamma}}_{j,\mathrm{I}}(a)$ via fitting (7). A covariate-adjusted estimator for $\tau_{j}(a,a^{\prime})$ can be obtained in the form of $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})=\widehat{\beta}_{j,\mathrm{I}}^{\mathrm{adj}}(a)-\widehat{\beta}_{j,\mathrm{I}}^{\mathrm{adj}}(a^{\prime})$, with $\widehat{\beta}_{j,\mathrm{I}}^{\mathrm{adj}}(a)=\overline{y}_{\cdot j\cdot}(a)-\overline{\bm{x}}_{\cdot j\cdot}^{c}(a)\widehat{\bm{\gamma}}_{j,\mathrm{I}}(a)$ where $\overline{\bm{x}}_{\cdot j\cdot}^{c}(a)=\sum_{i=1}^{I}\pi_{ij\cdot}G_{i}(a)\overline{\bm{X}}_{ij\cdot}^{c}/\allowbreak\pi_{\cdot j\cdot}(a)$ and $\overline{\bm{X}}_{ij\cdot}^{c}=\sum_{k=1}^{N_{ij}}\pi_{ijk}\bm{X}_{ijk}^{c}/\pi_{ij\cdot}$. The covariate-adjusted estimator $\widehat{\tau}_{j,\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})$ can be viewed as a multi-arm clustered extension of the regression estimator using a covariate-treatment indicator interaction model, as presented in Lin (2013).
@@
$\bm{E}_{\mathrm{I}}^{\mathrm{adj}}$ $=\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}\right)^{-1}\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{I}}\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{adj}}(\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}\right)$
$\qquad\times\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}})^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{adj}}$ is the vector of residuals from the WLS fit of (7), $\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}$ is of dimensions $N_{i\cdot}\times J(1+p_{x})(J+1)$ with $\bm{D}_{i,\mathrm{I}}^{\mathrm{adj}}=((\bm{D}_{i1,\mathrm{I}}^{\mathrm{adj}})^{\top}\cdots(\bm{D}_{iJ,\mathrm{I}}^{\mathrm{adj}})^{\top})^{\top}$ and the $k$-th row of $\bm{D}_{ij,\mathrm{I}}^{\mathrm{adj}}$ being $(\bm{G}_{i}(\mathcal{A})^{\top}\otimes\boldsymbol{\mathbbm{1}}_{ijk}^{\top},\bm{G}_{i}(\mathcal{A})^{\top}\otimes(\boldsymbol{\mathbbm{1}}_{ijk}^{\top}\otimes\bm{X}_{ijk}^{c}))$. The CR variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}\bm{Q}(a,a^{\prime})^{\top}$, where $\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{adj}}$ is the submatrix consisting of the first $J(J+1)$ columns and $J(J+1)$ rows of $\bm{E}_{\mathrm{I}}^{\mathrm{adj}}$. Let $\bm{\gamma}_{j,\mathrm{I}}(a)$ denote the WLS coefficient vectors obtained using individual-level data if the full set of potential outcomes is available. That is, regress $Y_{ijk}(a)$ on $(1,\bm{X}_{ijk}^{c})$ with weight $\pi_{ijk}$ assuming $Y_{ijk}(a)$ is available for each individual $k$ from cluster $i$ in period $j$. We further define the scaled cluster-period totals for non-centered and centered covariates as $\widetilde{\bm{X}}_{ij\cdot}=I\pi_{ij\cdot}\overline{\bm{X}}_{ij\cdot}$ and $\widetilde{\bm{X}}_{ij\cdot}^{c}=I\pi_{ij\cdot}\overline{\bm{X}}_{ij\cdot}^{c}$, respectively. The theoretical properties of $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})$ and $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}$ are given in Theorem 2.
@@
Define $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-\widetilde{\bm{X}}_{ij\cdot}^{c}\bm{\gamma}_{j,\mathrm{I}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\ldots,r_{iJ\cdot}(a))^{\top}$. Under Assumptions 1 and 2, and regularity conditions (C1) - (C3) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $\max_{i}\pi_{ij\cdot}=o(I^{-2/3})$ for $j=1,\ldots,J$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{adj}}(a,a^{\prime})\}=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.
@@
$\bm{E}_{\mathrm{I}}^{\mathrm{ancova}}$ $=\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}})^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}\right)^{-1}\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}})^{\top}\bm{\Pi}_{i,\mathrm{I}}\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{ancova}}(\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{ancova}})^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}\right)$
$\qquad\times\left(\sum_{i=1}^{I}(\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}})^{\top}\bm{\Pi}_{i,\mathrm{I}}\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}\right)^{-1},$
where $\widehat{\bm{e}}_{i,\mathrm{I}}^{\mathrm{ancova}}$ is the vector of residuals from the WLS fit of (8), $\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}$ is of dimensions $N_{i\cdot}\times J(J+1+p_{x})$ with $\bm{D}_{i,\mathrm{I}}^{\mathrm{ancova}}=((\bm{D}_{i1,\mathrm{I}}^{\mathrm{ancova}})^{\top}\cdots(\bm{D}_{iJ,\mathrm{I}}^{\mathrm{ancova}})^{\top})^{\top}$ and the $k$-th row of $\bm{D}_{ij,\mathrm{I}}^{\mathrm{ancova}}$ being $(\bm{G}_{i}(\mathcal{A})^{\top}\otimes\boldsymbol{\mathbbm{1}}_{ijk}^{\top},\boldsymbol{\mathbbm{1}}_{ijk}^{\top}\otimes\bm{X}_{ijk}^{c})$. The CR variance estimator for $\mathbf{Cov}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})\}$ is $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})\}=\bm{Q}(a,a^{\prime})\bm{E}_{\mathrm{I},[1\rightarrow J(J+1),1\rightarrow J(J+1)]}^{\mathrm{ancova}}\bm{Q}(a,a^{\prime})^{\top}$. Define $q(a)=I(a)/I$ as $I\rightarrow\infty$ for $a\in\mathcal{A}$. The theoretical properties of $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})$ and $\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})\}$ are given in the following Corollary 1.
@@
Define $r_{ij\cdot}(a)=\widetilde{\epsilon}_{ij\cdot}(a)-\widetilde{\bm{X}}_{ij\cdot}^{c}\sum_{a\in\mathcal{A}}q(a)\bm{\gamma}_{j,\mathrm{I}}(a)$ and $\bm{r}_{i\cdot}(a)=(r_{i1\cdot}(a),\allowbreak\ldots,r_{iJ\cdot}(a))^{\top}$. Under Assumptions 1 and 2, and regularity conditions (C1) - (C3) in Section S2 of the Online Supplement, for $a<a^{\prime}\in\mathcal{A}$, $\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})=\bm{\tau}(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$; if further $\max_{i}\pi_{ij\cdot}=o(I^{-2/3})$ for $j=1,\ldots,J$ and $\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime})\nrightarrow\bm{0}_{J\times J}$, then $I^{1/2}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})-\bm{\tau}(a,a^{\prime})\}\allowbreak\xrightarrow{d}\mathcal{N}(\bm{0}_{J},\bm{V}(\bm{r}_{i\cdot})(a,a^{\prime}))$, and $I\times\widehat{\mathbf{Cov}}_{\mathrm{CR}}\{\widehat{\bm{\tau}}_{\mathrm{I}}^{\mathrm{ancova}}(a,a^{\prime})\}=\bm{V}_{c}(\bm{r}_{i\cdot})(a,a^{\prime})+\bm{o}_{\mathbb{P}}(1)$.